# Vacuum transitions with the Gauss-Bonnet term in $D$ dimensions

### v2

@@ 3 Nucleation rates
According to (34), we know that $X\approx 1$ when $k_{+}$ is about zero and $k_{-}$ is finite. When we consider the down-tunnelling from near Minkowski space to anti-de Sitter space, which $k_{+}$ goes to zero while $k_{-}$ is finite, according to (32), the tunnelling exponent for $Min_{+}\rightarrow AdS_{-}$ goes as
@@ Appendix B Calculation of the boundary term in $D$-dimension
$K\equiv h^{ij}K_{ij}=(D-1)\frac{\rho^{\prime}}{\rho}.$ (48)
@@
$J_{ij}=-\frac{1}{3}(D-2)(D-3)\frac{\rho^{\prime 3}}{\rho^{3}}h_{ij},$ (49)
@@
$J\equiv h^{ij}J_{ij}=-\frac{1}{3}(D-1)(D-2)(D-3)\frac{\rho^{\prime 3}}{\rho^{3}}.$ (50)
@@
$\hat{G}_{ij}\equiv\hat{R}_{ij}-\frac{1}{2}h_{ij}\hat{R}=(D-2)\frac{1}{\rho^{2}}\delta^{ij}-\frac{1}{2}h^{ij}(D-1)(D-2)\frac{1}{\rho^{2}}.$ (51)
@@
$J-2\hat{G}_{ij}K_{ij}=(D-1)(D-2)(D-3)\left(\frac{\rho^{\prime}}{\rho^{3}}-\frac{1}{3}\frac{\rho^{\prime 3}}{\rho^{3}}\right).$ (52)
@@
$\int_{\Sigma}d^{D-1}x\sqrt{|h|}M^{D-2}_{pl}K=\int_{\Sigma}d^{D-1}xM^{D-2}_{pl}(D-1)\frac{\rho^{\prime}}{\rho},$ (53)
@@
$\int_{\Sigma}d^{D-1}x\sqrt{|h|}4\theta(J-2\hat{G}_{ij}K_{ij})=4\theta(D-1)(D-2)(D-3)\int_{\Sigma}d^{D-1}x\sqrt{|h|}\left(\frac{\rho^{\prime}}{\rho^{3}}-\frac{1}{3}\frac{\rho^{\prime 3}}{\rho^{3}}\right).$ (54)

### v3

@@ 3 Nucleation rates
When we consider the down-tunnelling from near Minkowski space to anti-de Sitter space, which $k_{+}$ goes to zero while $k_{-}$ is finite, according to (32), the tunnelling exponent for $Min_{+}\rightarrow AdS_{-}$ goes as
@@ Appendix B Calculation of the boundary term in $D$-dimension
$K\equiv\gamma^{ij}K_{ij}=(D-1)\frac{\rho^{\prime}}{\rho}.$ (48)
@@
$J_{ij}=-\frac{1}{3}(D-2)(D-3)\frac{\rho^{\prime 3}}{\rho^{3}}\gamma_{ij},$ (49)
@@
$J\equiv\gamma^{ij}J_{ij}=-\frac{1}{3}(D-1)(D-2)(D-3)\frac{\rho^{\prime 3}}{\rho^{3}}.$ (50)
@@
$\hat{G}_{ij}\equiv\hat{R}_{ij}-\frac{1}{2}\gamma_{ij}\hat{R}=(D-2)\frac{1}{\rho^{2}}\delta_{ij}-\frac{1}{2}\gamma_{ij}(D-1)(D-2)\frac{1}{\rho^{2}}.$ (51)
@@
$J-2\hat{G}_{ij}K^{ij}=(D-1)(D-2)(D-3)\left(\frac{\rho^{\prime}}{\rho^{3}}-\frac{1}{3}\frac{\rho^{\prime 3}}{\rho^{3}}\right).$ (52)
@@
$\int_{\Sigma}d^{D-1}x\sqrt{|\gamma|}M^{D-2}_{pl}K=\int_{\Sigma}d^{D-1}xM^{D-2}_{pl}(D-1)\frac{\rho^{\prime}}{\rho},$ (53)
@@
$\int_{\Sigma}d^{D-1}x\sqrt{|\gamma|}4\theta(J-2\hat{G}_{ij}K^{ij})=4\theta(D-1)(D-2)(D-3)\int_{\Sigma}d^{D-1}x\sqrt{|\gamma|}\left(\frac{\rho^{\prime}}{\rho^{3}}-\frac{1}{3}\frac{\rho^{\prime 3}}{\rho^{3}}\right).$ (54)